# Efficient and Flexible Method for Reducing Moderate-size Deep Neural Networks with Condensation

### v1

@@ Abstract
Neural networks have been extensively applied to a variety of tasks, achieving astounding results. Applying neural networks in the scientific field is an important research direction that is gaining increasing attention. In scientific applications, the scale of neural networks is generally moderate-size, mainly to ensure the speed of inference during application. Additionally, comparing neural networks to traditional algorithms in scientific applications is inevitable. These applications often require rapid computations, making the reduction of neural network sizes increasingly important. Existing work has found that the powerful capabilities of neural networks are primarily due to their non-linearity. Theoretical work has discovered that under strong non-linearity, neurons in the same layer tend to behave similarly, a phenomenon known as condensation. Condensation offers an opportunity to reduce the scale of neural networks to a smaller subnetwork with similar performance. In this article, we propose a condensation reduction algorithm to verify the feasibility of this idea in practical problems. Our reduction method can currently be applied to both fully connected networks and convolutional networks, achieving positive results. In complex combustion acceleration tasks, we reduced the size of the neural network to 41.7% of its original scale while maintaining prediction accuracy. In the CIFAR10 image classification task, we reduced the network size to 11.5% of the original scale, still maintaining a satisfactory validation accuracy. Our method can be applied to most trained neural networks, reducing computational pressure and improving inference speed.
@@ 1 Introduction
Neural networks have achieved globally astounding results, demonstrating exceptional performance across a range of tasks in scientific fields including biology, medicine, astronomy, environmental science, physics, and chemistry, etc [RNE+22, SK22, SBV20, SG23, ZLLT21, ZZB+21]. Serving as a novel numerical solution tool, neural networks have been effectively applied in solving partial differential equations within various domains [RPK19, BE21, LKA+20, MMOH20]. Neural networks can learn multiscale models from data by designing appropriate network structures and sampling methods, offering insights into these intricate systems [XYY+24].
@@
Currently, the scale of neural networks applied in the scientific field is generally moderate-size [RPK19]. Compared to the industrial and internet sectors, obtaining experimental data in the scientific field often incurs high costs, resulting in significantly fewer data samples available for neural network training [RPK19]. This limitation in sample size restricts the scale of networks that can be trained. In addition, training large-scale neural networks requires expensive computational resources. The computing conditions in academia generally lag behind those in the industrial sector, making it challenging to support the training and inference of overly large neural network models.
@@
Reducing the scale of neural networks applied in the scientific field holds significant importance [LSZ+18, WLW21]. Many traditional numerical algorithms, such as multigrid methods and fast Fourier transforms, exhibit quasi-linear or even linear time complexity, boasting extremely high computational efficiency [Wes95]. When applying neural networks to time-sensitive scientific computing tasks, it’s crucial to reduce the network size to enhance its inference speed [XYY+24]. Additionally, many scientific applications require deploying models in resource-constrained environments, such as embedded systems, mobile devices, and sensors [EMY+20, RSK+24]. The computational power of these deployment platforms often falls far short of the platforms used for the original training of the tasks. Therefore, it’s essential to reduce the network size as much as possible while maintaining the performance of the original network model, enabling efficient deployment in these contexts.
@@
Nonlinearity is the fundamental characteristic that endows neural networks with their powerful expressive, learning, and generative capacities [LXMZ21]. Non-linear activation functions grant neural networks the ability to learn complex patterns, enabling neurons to model complex nonlinear relationships and approximate any continuous function with arbitrary precision. The superposition of multiple layers of nonlinear transformations allows neural networks to construct highly nonlinear decision boundaries and learn complex nonlinear mappings between inputs and outputs. Theoretical findings indicate that in the presence of strong nonlinearity within neural networks, neurons in the same layer tend to exhibit a condensation phenomenon, known as condensation [LXMZ21]. Condensation refers to the gradual alignment of the direction of neurons’ parameter vectors during the training process, implying a uniform preference for the inputs from the previous layer. This phenomenon is prevalent across various neural network architectures, suggesting that condensed neurons function similarly, or are nearly equivalent to a single neuron. Condensation implies the presence of a large number of similar neurons in neural networks, indicating that the structural complexity of neural networks may not be as high as it appears. When the network reaches an extreme point, the condensation phenomenon becomes more pronounced. In this case, the embedding principle suggests that the extreme point reached by the network is actually an extreme point of one of its subnetworks [ZZLX21]. Therefore, in this article, we propose a reduction method to validate this idea.
@@
We are currently capable of reducing two major categories of neural networks: fully connected neural networks(FNN) and convolutional neural networks(CNN), achieving promising results. We selected two representative tasks for model reduction: fitting and classification. The fitting task involved neural network acceleration for combustion simulation, while the classification task focused on image classification on the CIFAR10 dataset, corresponding to fully connected neural networks and convolutional neural networks, respectively. In the acceleration of combustion simulation, we reduced the original model to 40.9% of its parameter count, maintaining consistency in both 0-dimensional and 1-dimensional combustion simulations with the original model, and observed virtually no difference in complex turbulent flame simulations. For the CIFAR10 image classification task, we reduced the model to 11.5% of its original parameter count, with the classification accuracy dropping only to 94% of the original accuracy.
@@ 2.2.1 Cosine similarity
When the cosine similarity between neurons is very close to $1$, their directions are almost identical.
@@ 3.1.2 Training Setup
wherein $m_{in}$ and $m_{in}$ are input dimension and output dimensions of the specific layer of the model. The initial learning rate is 0.0001, which decreases to 50% of its previous value every 1000 steps, with Adam as the optimizer. The starting batch size is 1024, which increases to 3.07 times its initial value every 1000 steps. The initial model undergoes training for 5000 steps, resulting in a training loss of 0.02172. Post-training, the model’s accuracy across the 0th, 1st, and 2nd dimensions is evaluated. The training and validation loss of the original model is shown in the figure 2(a).
@@ 3.1.3 Reduction Setup and Result
Due to the complex and time-consuming nature of evaluating neural network metrics in the combustion simulation task, a manual reduction approach is adopted. In the first reduction, the first layer of the original network is reduced, with a set condensation threshold of 0.9, resulting in a modified model architecture of 23-2205-1600-800-400-23. The initial learning rate is set at 0.0001, decreasing to 10% of its previous value every 2500 steps, with Adam as the optimizer. The starting batch size is 1024, increasing to 128 times its original size every 2500 steps. After training for 5000 steps, the training loss is 0.01592. This reduction in loss is considered indicative of the model’s performance being between that of the original model trained for 5000 steps and the original model trained for 10000 steps. This suggests that consolidating condensed neurons has a minimal impact on the neural network’s performance. Furthermore, it is observed from the loss graph that the loss of the reduced neural network rapidly decreases to a satisfactory level within a few training steps, similar to the case of perturbing the parameters of the original model. This indicates that the condensed neurons are approximately equivalent to a single neuron, a finding that is validated by subsequent reductions.The training and validation loss is shown in the figure 2(b).
@@
The third reduction targeted the second layer of the network, setting a condensation threshold of 0.999, resulting in a model architecture of 23-1105-1309-800-400-23. The condensation threshold was set so high because, after the first two reductions, the second layer of the neural network became highly condensed. A slightly lower condensation threshold would have led to a significant reduction in the network, potentially reducing the second layer to fewer than ten neurons and greatly impacting model performance. The phenomenon where reducing one layer of the neural network causes a higher degree of condensation in the subsequent layer is interesting. In practice, the network parameter matrices for the first and second layers, which we use for reduction, can be understood as the input and output parameter matrices of the first layer of the neural network, respectively. They are inherently related, and reducing the first layer of the network decreases the number of rows in the input parameter matrix while also decreasing the number of columns in the output parameter matrix. The relationship between them will be explored further in future research. As shown in the figure 3, it can be clearly seen that the reduction of the first layer of the neural network can immediately change the condensation state of the second layer of the neural network. After the first and second reductions, when examining the cosine similarity matrix of the second layer of the neural network in the untrained state, we can see that the previously inconspicuous condensation in the second layer appeared after the first reduction, and a strong condensation phenomenon emerged after the second reduction, making the second layer neural network capable of extremely strong condensation even at a high condensation threshold of 0.9. Since the previous model had already been trained for 15000 steps and reached a vicinity of a good extreme point, a smaller learning rate was used for this training phase. The initial learning rate was adjusted to 4e-5, with all other settings remaining the same as in the first reduction. After training for 5000 steps, the training loss was recorded at 0.01360, showing almost no change from the second reduction.The training and validation loss is shown in the figure 2(d). The results of the two-dimensional turbulence simulation of the reduced model are shown in the figure 4, which is almost identical to the original model.
@@
Figure 3: Cosine similarity matrixs of 2nd layer on different stages.The more distinct the blocks in the cosine matrix, the stronger the condensation of the neural network. From left to right, the images correspond to the original fully connected neural network model through to the model after the third reduction. ”Epoch 0” indicates that the model has just been reduced and has not yet been trained. ”Epoch 5000” represents the fully connected neural network model upon completion of training.
@@ 3.2.1 Background on task
MobileNetV2 is a convolutional neural network model that adopts depthwise separable convolutions and is dedicated to mobile device deployment [SHZ+18]. Its main network structure is shown in the tabel 1. For detailed information about MobileNetV2, please refer to the original text [SHZ+18]. It concatenates 16 modules with the same structural characteristics, and the structure is shown in the figure 5. The reasons for using MobileNetV2 for condensation reduction are as follows. First, MobileNetV2 targets mobile deployment scenarios, which is also an important application scenario for model reduction. MobileNetV2 itself contains a small number of parameters, and further reduction of MobileNetV2 can reflect the reduction capability and practical application value of the condensation reduction method. Second, MobileNetV2 has 16 modules with the same structural characteristics, which is very suitable for using automatic condensation reduction methods for layer-by-layer reduction. Finally, MobileNetV2 mainly adopts depthwise separable convolutions, which will further extend the definition of condensation and enrich the application fields of condensation reduction.
@@ 3.2.2 Training and Reduction Setup
wherein $m_{in}$ and $m_{in}$ are input dimension and output dimensions of the specific layer of the model. The initial learning rate is set at 0.01, with a cosine annealing strategy employed for the first 400 steps, featuring a cycle length of 200 and a minimum learning rate of 0, using SGD as the optimizer. After 400 steps, the learning rate shifts to 0.001. Subsequently, every 1000 steps, the learning rate alternates between 0.001 and 0.0001. The main idea behind this learning rate strategy is to first find a suitable parameter region for the neural network using the cosine annealing strategy, and then refine the details of the neural network parameters with Adam. Throughout the training process, the batch size is maintained at 128.
@@
wherein $t_{reduction}$ represents the current time of reductions, $T_{lr}$ is set to 200, $lr_{min}$ is set to 0.0001 and $lr_{max}$ is set to 0.01. The model initiates main reduction process when the accuracy on the validation set reaches 0.88. Once the model achieves this main reduction criterion, it automatically begins layer-by-layer reduction. The criterion for layer-by-layer reduction is set at an accuracy of 0.84 on the validation set. After initiating layer-by-layer reduction, whenever the model meets this criterion, it automatically reduces the corresponding neural network layer. After each reduction, both the main reduction and layer-by-layer reduction criterion are automatically adjusted in the following manner:

### v2

@@ Abstract
Neural networks have been extensively applied to a variety of tasks, achieving astounding results. Applying neural networks in the scientific field is an important research direction that is gaining increasing attention. In scientific applications, the scale of neural networks is generally moderate size, mainly to ensure the speed of inference during application. Additionally, comparing neural networks to traditional algorithms in scientific applications is inevitable. These applications often require rapid computations, making the reduction in neural network sizes increasingly important. Existing work has found that the powerful capabilities of neural networks are primarily due to their nonlinearity. Theoretical work has discovered that under strong nonlinearity, neurons in the same layer tend to behave similarly, a phenomenon known as condensation. Condensation offers an opportunity to reduce the scale of neural networks to a smaller subnetwork with a similar performance. In this article, we propose a condensation reduction method to verify the feasibility of this idea in practical problems, thereby validating existing theories. Our reduction method can currently be applied to both fully connected networks and convolutional networks, achieving positive results. In complex combustion acceleration tasks, we reduced the size of the neural network to 41.7% of its original scale while maintaining prediction accuracy. In the CIFAR10 image classification task, we reduced the network size to 11.5% of the original scale, still maintaining a satisfactory validation accuracy. Our method can be applied to most trained neural networks, reducing computational pressure and improving inference speed.
@@ 1 Introduction
Neural networks have achieved globally astounding results, demonstrating an exceptional performance across a range of tasks in scientific fields including biology, medicine, astronomy, environmental science, physics, chemistry, etc. [RNE+22, SK22, SBV20, SG23, ZLLT21, ZZB+21]. Serving as a novel numerical solution tool, neural networks have been effectively applied in solving partial differential equations within various domains [RPK19, BE21, LKA+20, MMOH20]. Neural networks can learn multiscale models from data by designing appropriate network structures and sampling methods, offering insights into these intricate systems [XYY+24].
@@
Currently, the scale of neural networks applied in the scientific field is generally moderate size [RPK19]. Compared to the industrial and internet sectors, obtaining experimental data in the scientific field often incurs high costs, resulting in significantly fewer data samples available for neural network training [RPK19]. This limitation in sample size restricts the scale of networks that can be trained. In addition, training large-scale neural networks requires expensive computational resources. The computing conditions in academia generally lag behind those in the industrial sector, making it challenging to support the training and inference of overly large neural network models.
@@
Reducing the scale of neural networks applied in the scientific field holds significant importance [LSZ+18, WLW21]. Many traditional numerical algorithms, such as multigrid methods and fast Fourier transforms, exhibit quasi-linear or even linear time complexity, boasting an extremely high computational efficiency [Wes95]. When applying neural networks to time-sensitive scientific computing tasks, it is crucial to reduce the network size to enhance its inference speed [XYY+24]. Additionally, many scientific applications require deploying models in resource-constrained environments, such as embedded systems, mobile devices, and sensors [EMY+20, RSK+24]. The computational power of these deployment platforms often falls far short of the platforms used for the original training of the tasks. Therefore, it is essential to reduce the network size as much as possible while maintaining the performance of the original network model, enabling efficient deployment in these contexts.
@@
Nonlinearity is the fundamental characteristic that endows neural networks with their powerful expressive, learning, and generative capacities [LXMZ21]. Nonlinear activation functions grant neural networks the ability to learn complex patterns, enabling neurons to model complex nonlinear relationships and approximate any continuous function with arbitrary precision. The superposition of multiple layers of nonlinear transformations allows neural networks to construct highly nonlinear decision boundaries and learn complex nonlinear mappings between inputs and outputs. Theoretical findings indicate that in the presence of strong nonlinearity within neural networks, neurons in the same layer tend to exhibit a condensation phenomenon, known as condensation [LXMZ21]. Condensation refers to the gradual alignment of the direction of neurons’ parameter vectors during the training process, implying a uniform preference for the inputs from the previous layer. This phenomenon is prevalent across various neural network architectures, suggesting that condensed neurons function similarly, or are nearly equivalent to a single neuron. Condensation implies the presence of a large number of similar neurons in neural networks, indicating that the structural complexity of neural networks may not be as high as it appears. When the network reaches an extreme point, the condensation phenomenon becomes more pronounced. In this case, the embedding principle suggests that the extreme point reached by the network is actually an extreme point of one of its subnetworks [ZZLX21].
Based on the above idea, we propose the condensation reduction method to validate existing theories. The main contribution of this paper is the successful validation of the universality of the condensation phenomenon and the applicability of the embedding principle in practical applications through the condensation reduction method. The use of the condensation reduction method requires a pretrained network, which could be considered a drawback. However, this limitation might be quite common, as seen in the work related to the “Lottery Ticket Hypothesis” [FC18].
From the perspective of reduction algorithms, our method can be approximated as a pruning method. The key difference is that our approach merges branches instead of merely removing redundant ones. The work by Liu et al. [LZS+21] proposed a network condensation method, but their definition of condensation differs from ours. In their work, the term “condense” is primarily used as a verb to denote operations such as simplifying, refining, and compressing the network. However, in this study, “condensation” is defined as a significant phenomenon commonly observed during neural network training and can be directly observed using cosine similarity. Most pruning methods measure the importance of parameters based on certain metrics and remove the less important ones [CZS23], which can be either structured or unstructured. These metrics usually include parameter norms, impact on loss, and sensitivity [HP88, LDS89, YYY+19].
In contrast, our condensation reduction method is a structured network reduction approach. Unlike structured pruning, we do not delete neurons or parameters based on metrics. Instead, we merge neurons based on their similarity (degree of condensation). Merging neurons involves deleting redundant neurons and modifying the parameters of the retained neurons, ensuring that the merged neurons have similar expressive capabilities as before. In the following sections, we prove the consistency of the neural network’s output before and after the merger under strict conditions.
@@
We are currently capable of reducing two major categories of neural networks: fully connected neural networks (FNNs) and convolutional neural networks (CNNs), achieving promising results. We selected two representative tasks for model reduction: fitting and classification. The fitting task involved neural network acceleration for combustion simulation, while the classification task focused on image classification on the CIFAR10 dataset, corresponding to fully connected neural networks and convolutional neural networks, respectively. In the acceleration of the combustion simulation, we reduced the original model to 40.9% of its parameter count, maintaining consistency in both zero-dimensional and one-dimensional combustion simulations with the original model, and observed virtually no difference in complex turbulent flame simulations. For the CIFAR10 image classification task, we reduced the model to 11.5% of its original parameter count, with the classification accuracy dropping only to 94% of the original accuracy.
@@ 2.2.1 Cosine similarity
When considering neurons as vectors, if the cosine similarity between them is very close to 1, their directions are very similar, indicating that these two neurons are highly similar. The calculation of the cosine similarity is straightforward. Furthermore, in the following section, we will demonstrate that when the cosine similarity is exactly $1$, the functions expressed by the original network and the reduced network are consistent.
@@ 3.1.2 Training Setup
wherein $m_{in}$ and $m_{in}$ are input dimension and output dimensions of the specific layer of the model. The initial learning rate is set to 0.0001 and is reduced to 50% of its original value every 1000 steps to prevent the learning rate from becoming too large. The entire training process uses the Adam optimizer, sets $betas=(0.9,0.999)$, $eps=1\times 10^{-8}$, $weight\leavevmode\nobreak\ decay=0$. The starting batch size is set to 1024 and is increased by a factor of 3.07 every 1000 steps to gradually accelerate training and enhance the model’s generalization performance. The initial model undergoes training for 5000 steps, resulting in a training loss of 0.02172. Post-training, the model’s accuracy across the 0th, 1st, and 2nd dimension is evaluated. The training and validation loss of the original model is shown in Figure 2a.
@@ 3.1.3 Reduction Setup and Result
Due to the complex and time-consuming nature of evaluating neural network metrics in the combustion simulation task, a manual reduction approach is adopted. Please refer to Table 1 for a comparison of the reduction processes. In the first reduction, the first layer of the original network is reduced, with a set condensation threshold of 0.9, resulting in a modified model architecture of 23-2205-1600-800-400-23. The initial learning rate is set at 0.0001, decreasing to 10% of its previous value every 2500 steps, with Adam as the optimizer. The starting batch size is 1024, increasing to 128 times its original size every 2500 steps. After training for 5000 steps, the training loss is 0.01592. This reduction in loss is considered indicative of the model’s performance being between that of the original model trained for 5000 steps and the original model trained for 10000 steps. This suggests that consolidating condensed neurons has a minimal impact on the neural network’s performance. Furthermore, it is observed from the loss graph that the loss of the reduced neural network rapidly decreases to a satisfactory level within a few training steps, similar to the case of perturbing the parameters of the original model. This indicates that the condensed neurons are approximately equivalent to a single neuron, a finding that is validated by subsequent reductions.The training and validation loss is shown in the figure 2(b).
@@
The third reduction targeted the second layer of the network, setting a condensation threshold of 0.999, resulting in a model architecture of 23-1105-1309-800-400-23. The condensation threshold was set so high because, after the first two reductions, the second layer of the neural network became highly condensed. A slightly lower condensation threshold would have led to a significant reduction in the network, potentially reducing the second layer to fewer than ten neurons and greatly impacting model performance. The phenomenon where reducing one layer of the neural network causes a higher degree of condensation in the subsequent layer is interesting. In practice, the network parameter matrices for the first and second layers, which we use for reduction, can be understood as the input and output parameter matrices of the first layer of the neural network, respectively. They are inherently related, and reducing the first layer of the network decreases the number of rows in the input parameter matrix while also decreasing the number of columns in the output parameter matrix. The relationship between them will be explored further in future research. As shown in the figure 3, it can be clearly seen that the reduction of the first layer of the neural network can immediately change the condensation state of the second layer of the neural network. After the first and second reductions, when examining the cosine similarity matrix of the second layer of the neural network in the untrained state, we can see that the previously inconspicuous condensation in the second layer appeared after the first reduction, and a strong condensation phenomenon emerged after the second reduction, making the second layer neural network capable of extremely strong condensation even at a high condensation threshold of 0.9. Since the previous model had already been trained for 15000 steps and reached a vicinity of a good extreme point, a smaller learning rate was used for this training phase. The initial learning rate was adjusted to 4e-5, with all other settings remaining the same as in the first reduction. After training for 5000 steps, the training loss was recorded at 0.01360, showing almost no change from the second reduction.The training and validation loss is shown in the figure 2(d). The results of the two-dimensional turbulence simulation of the reduced model are shown in the figure 4, which is almost identical to the original model. According to existing theories, we know that training a large network and gradually reducing its size can yield better results compared to directly training a small network. We retrained a small network with the same size as the reduced network, using the same training settings as previously described. The directly trained small network even failed to converge when predicting two-dimensional turbulent flames.
@@
Figure 3: Cosine similarity matrixs of 2nd layer on different stages. In the matrix heatmap, the element in the $i$-th row and $j$-th column represents the cosine similarity between the $i$-th neuron and the $j$-th neuron. The more distinct the blocks in the cosine matrix, the stronger the condensation of the neural network. From left to right, the images correspond to the original fully connected neural network model through to the model after the third reduction. ”Epoch 0” indicates that the model has just been reduced and has not yet been trained. ”Epoch 5000” represents the fully connected neural network model upon completion of training.
@@
Table 1: Comparison table of reduction process of FNN. “Parameter” represents the number of parameters in the neural network. Reduction rate represents the ratio of the current network parameter count to the original network parameter count. A lower reduction rate indicates a smaller network relative to the original, demonstrating a more effective reduction.
Model Structure Validation Loss Parameter reduction ratio
Original 23-3200-1600-800-400-23 0.02276 6,802,800 100%
1st reduced 23-2205-1600-800-400-23 0.01694 5,187,915 76.26%
2nd reduced 23-1105-1600-800-400-23 0.01454 3,402,615 50.01%
3rd reduced 23-1105-1309-800-400-23 0.01434 2,848,260 41.87%
@@ 3.2.1 Background on task
MobileNetV2 is a convolutional neural network model that adopts depthwise separable convolutions and is dedicated to mobile device deployment [SHZ+18]. Its main network structure is shown in the tabel 2. For detailed information about MobileNetV2, please refer to the original text [SHZ+18]. It concatenates 16 modules with the same structural characteristics, and the structure is shown in the figure 5. The reasons for using MobileNetV2 for condensation reduction are as follows. First, MobileNetV2 targets mobile deployment scenarios, which is also an important application scenario for model reduction. MobileNetV2 itself contains a small number of parameters, and further reduction of MobileNetV2 can reflect the reduction capability and practical application value of the condensation reduction method. Second, MobileNetV2 has 16 modules with the same structural characteristics, which is very suitable for using automatic condensation reduction methods for layer-by-layer reduction. Finally, MobileNetV2 mainly adopts depthwise separable convolutions, which will further extend the definition of condensation and enrich the application fields of condensation reduction.
@@ 3.2.2 Training and Reduction Setup
wherein $m_{in}$ and $m_{out}$ the are input dimension and output dimension of the specific layer of the model.
The initial learning rate is set to 0.01, and a cosine annealing strategy with a period length of 200 and a minimum learning rate of 0 is used for the first 400 steps. The optimizer during this phase is SGD with $momentum=0.9$. This part of the training can be seen as the process of quickly finding a good critical point for the model. After 400 steps, the learning rate changes to 0.001, and the optimizer switches to Adam, sets $betas=(0.9,0.999)$, $eps=1\times 10^{-8}$, $weight\leavevmode\nobreak\ decay=0$. Subsequently, every 1000 steps, the learning rate alternates between 0.001 and 0.0001. The main idea behind this learning rate strategy is to first use the cosine annealing strategy to find a suitable parameter region for the neural network and then use the Adam optimizer to refine the network parameters.
In our multiple experiments, we observed a noticeable improvement in test accuracy after switching to Adam. The accuracy obtained using this mixed strategy was higher compared to using Adam throughout the entire training process. Throughout the entire training process, the batch size remained at 128.
@@
wherein $t_{reduction}$ represents the current time of reductions, $T_{lr}$ is set to 200, $lr_{min}$ is set to 0.0001 and $lr_{max}$ is set to 0.01. The reason for this approach is that each reduction can be considered as the model reaching a critical point that meets our requirements. At this point, the learning rate should be appropriately lowered to further enhance the model’s performance. The model initiates main reduction process when the accuracy on the validation set reaches 0.88. Once the model achieves this main reduction criterion, it automatically begins layer-by-layer reduction. The criterion for layer-by-layer reduction is set at an accuracy of 0.84 on the validation set. After initiating layer-by-layer reduction, whenever the model meets this criterion, it automatically reduces the corresponding neural network layer. After each reduction, both the main reduction and layer-by-layer reduction criterion are automatically adjusted in the following manner: